Assume your organization has 100 patients analyzed in the following manner: 15 Medicare patients, who pay$2,000 per diagnosis 25 Medicaid patients, who pay $1, 800 per diagnosis 20 managed care patients, who pay charges minus a 20% discount 10 managed care patients, who pay charges minus a 25% discount 10 private insurance patients, who pay charges 10 charity care patients, who pay nothing 10 bad debt patients, who pay nothing Your organization's average cost per patient is $2,000. Calculate the charge necessary to recover your cost How would you put this an excel
Answer #1

Answer:

TOTAL NO.OF PATIENTS=100
COST PER PATIENT=$2,000
TOTAL COST TO BE RECOVERED=$2,000*100=$200,000
LET THE CHARGE FROM EACH PATIENT BE X
PAYMENT RECEIVED:
NO.OF PATIENTS AMOUNT($) Calculation
15 30,000 (2000*15)
25 45,000 (1800*25)
20 20*X-20%=16X
10 10*X-25%=7.5X
10 X
10 -
10 -
TOTAL 75,000+24.5X
THEREFOR $75,000+24.5X=$200,000
24.5X=125,000
X=$5,102.04
THE CHARGE NECESSARY TO RECOVER THE COST IS $ 5,102.04